# Seamless Monitoring of Stress Levels Leveraging a Universal Model for Time Sequences

### v1

@@ 3.3 Fine-tuning UniTS for Abnormal Emotion Detection
UniTS was trained on a large, diverse dataset using a unified masked reconstruction scheme. This allows the model to learn general representations of temporal sequences and their underlying patterns, enhancing its generative and predictive task capabilities. Recall that UniTS is a foundational model in time series, and its training is executed on multiple tasks.
UniTS enables two different adaptation strategies for new datasets: i.e., whole model fine-tuning and prompt learning. The model weights are frozen in prompt learning, and only the prompt tokens are updated. According to [12], this performs similarly to fine-tuning at a lower computational cost. Nevertheless, because we want to demonstrate the capability of HR and HRV signals only to detect abnormal emotion detection via non-invasive wearable devices, we preprocess the datasets to obtain features at a low-frequency rate – see Sec. 4.2. This makes fine-tuning UniTS more “affordable” since we drastically reduce the sampling rate of the signals present in the dataset (e.g., from a sample of ECG of 700Hz to a transformed sample of HR/HRV of 0.1Hz). We follow [6] to reduce the sampling rate – this is also supported in Figure 3 – since we want to emulate the realness of how vital signals are measured with smartwatches while maintaining effective battery levels. Finally, note that we use 20 epochs and an exponential learning rate decay to fine-tune the model.
4 Experiments
@@ 4.1 Experimental Setup
Compared methods. We compare supervised UniTS and fine-tuned UniTS with various SoTA models for multivariate time-series anomaly detection. The models compared include TranAD[37], LSTM-NDT[19], CAE-M[41], DAGMM[45], USAD[4], OmniAnomaly[35], MTAD-GAT[42], GDN[10], MAD-GAN[25], MSCRED[39], HypAD[11], and TadGAN[13]. For these models, we adopted the hyperparameters as presented in their original papers and conducted a learning rate grid search using 5-fold cross-validation to ensure optimal performance.
Training UniTS. For fine-tuning UniTS, we used a window size of 5 and an embedding dimension of 128. These hyperparameters were selected based on an ablation study on the model. The datasets were divided into 80% training data and 20% validation data. All models were tested every 5 epochs since different convergence times may occur, and the best result for each fold was recorded and averaged to determine the optimal learning rate.
@@
Fair comparisons and threshold settings. We systematically preprocessed our datasets to ensure consistency and comparability across different sources. From all datasets, we extracted Heart Rate (HR) and Heart Rate Variability (HRV) every 10 seconds using a sliding window approach based on the preceding 60 seconds of signal data from ECG and, when available, Blood Volume Pulse (BVP), which measure the volumetric change in blood circulation through the microvascular bed of tissue. BVP signals can be extracted non-invasively using photoplethysmography (PPG), a technique commonly integrated into smartwatches and other wearable devices. This capability is instrumental in stress detection and management, making wearable devices valuable in clinical and everyday settings.
Lastly, we use a threshold of the 3rd and 97th percentile to indicate anomalous data points: i.e., everything that is below the 3rd or above the 97th percentile is labeled as 1; otherwise, it is 0. Percentiles are calculated on the reconstruction errors.
@@ 4.2 Datasets
We use three widely adopted datasets for emotion detection:
DREAMER [20] is designed for emotion recognition via low-cost devices. It consists of EEG and ECG readings recorded from 23 participants. ECG signals were captured at a sampling rate of 256 Hz. Movie clips were used to stimulate different emotions in participants, and at the end of each clip, participants scored their self-assessment on a 5-point arousal, valence, and dominance scale. Here, we treat states with a valence of 1 as abnormal and the rest as normal.
MAHNOB-HCI [34] was collected using the Biosemi active II system with active electrodes to capture physiological signals, among others the ECG and EEG, from 27 participants at a sampling rate of 256 Hz. Participants were asked to record their emotional status after each trial, consisting of multimedia content stimulation. Here, we treat the states with a valence of 1 as abnormal and the rest of the emotional states as normal. For our experiments, we selected subsets of the dataset with recordings of at least 120 seconds and excluded those with incomplete tracks due to recording errors.
WESAD [33] has ECG and BVP signals from 15 participants at a rate of 700Hz extracted via a RespiBAN Professional 2 sensor worn on the chest. It contains four different emotional states. Here, we treat the stress state as abnormal and the rest as normal. The stress condition is obtained by exposing the subjects to the Tier Social Stress Test, which consists of public speaking and a mental arithmetic task.
Table 2: Quantitative experiments of UniTS vs SoTA methods on three benchmarking datasets.
(a) Comparison of different methods using HR and HRV extracted from ECG and BVP. Bold-faced values illustrate the best; underlined ones are the second best.
@@
LSTM-NDT 0.313 0.375 0.785 0.772 0.561 $\pm$ .218
OmniAnomaly 0.599 0.547 0.767 0.650 0.641 $\pm$ .081
CAE-M 0.658 0.530 0.801 0.590 0.645 $\pm$ .101
HypAD 0.650 0.643 0.815 0.567 0.669 $\pm$ .090
MTAD-GAT 0.536 0.742 0.857 0.719 0.714 $\pm$ .115
TranAD 0.617 0.623 0.837 0.804 0.720 $\pm$ .101
TadGAN 0.590 0.691 0.864 0.743 0.722 $\pm$ .099
GDN 0.713 0.580 0.858 0.802 0.738 $\pm$ .105
DAGMM 0.743 0.647 0.831 0.773 0.749 $\pm$ .067
USAD 0.730 0.660 0.830 0.797 0.754 $\pm$ .065
MAD-GAN 0.706 0.743 0.839 0.787 0.769 $\pm$ .050
MSCRED 0.675 0.824 0.876 0.775 0.788 $\pm$ .074
UniTS 0.869 0.878 0.834 0.856 0.859 $\pm$ .019
(b) P-values produced by the Dunn post-hoc test, where the control model is UniTS against the top-performing SoTA methods. We use a p-value of $5\times 10^{-2}$ to reject the null hypothesis.
UniTS
MSCRED $2.813\times 10^{-2}$
MAD-GAN $4.795\times 10^{-3}$
USAD $6.839\times 10^{-3}$
DAGMM $1.136\times 10^{-3}$
GDN $5.146\times 10^{-3}$
@@ 4.3 Quantitative Results
Overall, UniTS exhibits remarkably stable performance and higher robustness across all datasets and systems (std of only .019.019.019.019), resulting from being a universal model pre-trained on many diverse signals and tasks.
UniTS achieves performances from lightweight devices comparable to SoTA systems using more invasive and potentially accurate sensors. The only dataset collecting signals from ECG and smartwatch (BVP) is WESAD. Table 2(a) (columns 4 and 5) clearly shows that performance worsens with BPV for all systems except UniTS, which maintains similar F1 values for the ECG and BPV versions. This is a valuable result since, as discussed in Section 1, lightweight devices, such as bracelets and smartwatches, are less expensive and invasive, enabling seamless monitoring. Using non-invasive, wearable technology without sacrificing performance can expand stress-related research and applications. Future studies can integrate these devices in various settings to gather continuous stress data, leading to personalized stress management programs, early intervention systems for mental health, and better well-being monitoring.
Lower sampling rates on lightweight devices makes performance plummet by $\mathbf{\sim 66\%}$. Figure 3 shows the importance of the sampling rate on all datasets on ECG and BVP versions. Here, we illustrate the trend of average F1 scores of UniTS when the sampling interval increases from 10 to 60 seconds. The vital signals (i.e., HR and HRV) are sampled every $k$ seconds to build the multivariate time series taken in input, with varying $k$. As expected, the longer the sampling interval, the less responsive the model becomes, resulting in a degradation of results. This phenomenon is even more visible when we extract the HR and HRV from BVP, having a drop of $-66.12\%$ passing from a 10 to a 60-second sampling rate.
Figure 3: Performance degradation of F1 score across different sampling intervals using the UniTS on all the datasets.
This discrepancy highlights a critical consideration in the evolving landscape of wearable technology and continuous monitoring systems. While Malik et al. [29] provided foundational guidelines for HRV measurement, recommending high sampling rates of at least 200Hz for precise R-R interval [^8] https://www.sciencedirect.com/topics/nursing-and-health-professions/rr-interval detection from ECG and 2-to-5-minutes for BVP, our empirical findings suggest that for real-time HRV analysis, using BVP signals, the sampling intervals must be significantly shorter to maintain model performance. We suggest a re-evaluation of the balance between device convenience and performance. For applications requiring high precision, such as medical diagnostics or performance monitoring in athletes, adhering to more frequent sampling intervals appears imperative. Therefore, our study advocates for a nuanced approach, where the selection of sampling intervals is tailored to the specific requirements of the use case, ensuring that the integrity of physiological data is not compromised by the constraints of the monitoring device.
@@
4.4 Qualitative Results
@@
Casting the problem as one of anomaly detection enhances explainability. Anomaly detection begins by establishing a model of ”normal" behavior based on the monitored signals, and next, it detects deviations from this normality. This helps clinicians understand the baseline from which deviations are identified and aligns better with medical professionals’ needs and practical constraints. First, clinicians are usually more interested in identifying and investigating abnormalities rather than confirming the norm. Anomaly detection systems align with doctors’ diagnostic workflow by focusing on detecting deviations from the norm. Secondly, these models provide significant benefits for precision medicine and personalization due to their ability to identify unique patient profiles and adapt to new data continuously [30]. Furthermore, in the case of multiple monitored signals, the model should highlight which ones are more indicative of stress [3]. In Figure 4, we aim to support doctors in identifying which signals contributed the most in detecting anomalies. Here, we depict the contribution probability based on the reconstruction error of each signal – i.e., HR (up) and HRV (middle) – in determining anomalies. The more opaque the shade of each bar in the plot, the more likely an anomaly will occur. Notice how the opacity of the probabilities increases when there is an abrupt trend change in one (or both) of the signals, which aligns with the red areas depicting the detection of an anomaly. One can think of the transparency of each bar as a visual uncertainty of the model. Therefore, the more opaque the bars are, the more certain the model is that there is an anomaly in that particular time frame. Meanwhile, when the bars are transparent, they may indicate two things: either the model is uncertain, or the data points therein are normal. To this end, we integrate the contribution probability with detecting anomalies in the raw signals. Additionally, we invite the reader to notice the different transparency levels within the time frame where an anomaly is detected. This phenomenon is expected as not all types of anomalies (gradual vs. abrupt) are detected with the same confidence. Interestingly, in this scenario, badly reconstructing HR pushes the model to be more confident about an anomaly occurring – see the middle portion of the third row in Figure 4. Contrarily, even if UniTS cannot effectively reconstruct HRV, it does not lead to detecting anomalies – e.g., notice the more transparent tail-ends of the probabilities plot. By exploiting these types of explicit and inherently interpretable visualizations – i.e., the contribution probability is estimated via the reconstruction error – doctors can interpret which signal guides the prediction of an anomaly during inference and intervene accordingly. Finally, this type of analysis can be used to debug if the monitoring system functions correctly. For instance, if the contribution probability bars are opaque, yet the system does not report an anomaly, there might be something wrong with the current monitoring setup.
5 Conclusion
This is the first study to employ a universal time series model, UniTS, to support the seamless monitoring of physiological signals. We have demonstrated that UniTS significantly outperforms state-of-the-art machine learning models for stress detection. Furthermore, it performs similarly when using invasive monitoring devices, such as the ECG, and lightweight, such as a smartwatch or bracelet. This equivalence is valid as long as the sampling frequency is sufficiently high but still well within the reach of existing devices, even less expensive ones. Monitoring patients’ disease progression via lightweight devices has numerous advantages, including increased patient comfort and compliance, continuous and real-time monitoring, reduced healthcare costs, increased patient independence, personalized care, improved remote management of patients, and better health outcomes through early intervention.

### v2

@@ 3.3 Fine-tuning UniTS for Abnormal Emotion Detection
UniTS was trained on a large, diverse dataset using a unified masked reconstruction scheme. This allows the model to learn general representations of temporal sequences and their underlying patterns, enhancing its generative and predictive task capabilities. Recall that UniTS is a foundational model in time series, and its training is executed on multiple tasks. In UniTS, whole model fine-tuning and prompt-learning represent two distinct paradigms for adapting the foundational model to new datasets and specific tasks, each with unique advantages and operational mechanisms.
Fine-tuning vs. Prompt-learning
Fine-tuning involves updating the entire UniTS model parameters using the target dataset. This approach leverages the pre-trained representations learned during the initial training phase and adapts them to the nuances of the new data. Specifically, fine-tuning adjusts the weights of both the sequence and variable attention mechanisms and the dynamic linear operators within the unified network backbone. This comprehensive adaptation allows the model to specialize its understanding and prediction capabilities to better fit the characteristics of the new dataset. Prompt learning, on the other hand, maintains the integrity of the pre-trained UniTS model by keeping its weights frozen and introducing a set of learnable prompt tokens. These prompt tokens act as task-specific adapters that guide the model’s processing without altering its foundational parameters. Only the embeddings of these prompt tokens are updated during prompt learning based on the target task, effectively steering the model to generate appropriate outputs for the new dataset. Table 2 illustrates these paradigms’ (dis)advantages.
Table 2: Comparison of Fine-tuning and Prompt-Learning in UniTS.
Fine-tuning Prompt-Learning
Advantages
 Full task-specific optimization.
 Enhanced performance for diverse datasets.
  Efficient with frozen model weights.
  Rapid adaptation to new tasks.
  Reduced risk of overfitting.
Limitations
  Expensive to update model parameters.
  Overfitting on small datasets.
  Lower performance vs. fine-tuning.
  Effectiveness depends on prompt design.
According to Gao et al. [14], prompt learning in UniTS achieves performance comparable to fine-tuning while incurring lower computational costs. This parity is particularly advantageous in applications where resource constraints are a concern or when deploying the model across multiple tasks simultaneously. However, fine-tuning remains the preferred approach when maximum performance on a specific task is paramount and computational resources are available.
In the context of our study, where the objective is to utilize HR and HRV signals for abnormal emotion detection via non-invasive wearable devices, the choice between fine-tuning and prompt learning is influenced by practical considerations. By preprocessing the datasets to obtain features at a low-frequency rate (e.g., transforming ECG samples from 700Hz to HR/HRV at 0.1Hz), we significantly reduce the computational burden of fine-tuning UniTS. This reduction aligns with our goal to emulate real-world scenarios where vital signals are measured with smartwatches, balancing effective battery usage with reliable signal acquisition [7]. Consequently, fine-tuning becomes a more feasible and “affordable” strategy in our application, enabling the model to adapt effectively without compromising efficiency. We follow [7] to reduce the sampling rate – this is also supported in Figure 5 of subsection 4.4. Finally, we use 20 epochs and an exponential learning rate decay to finetune the model.
3.4 Adding explainability features
Once the UniTS model is finetuned for anomaly detection, we leverage the output of the Tower module during inference (as illustrated in Figure 1) to serve as input for a Large Language Model (LLM), such as GPT-4o [1] or LLaMa-2 [40], to generate explanations for the detected anomalies. Specifically, the anomaly score generated by UniTS represents the likelihood of an anomaly, calculated as the ratio of detected anomaly energy to a set threshold. This score indicates the intensity of the anomaly, with values below 1 denoting normal samples and values above 1 indicating an anomaly. This score is then used to prompt the LLM to generate interpretive explanations for domain experts.
The LLM can be prompted to contextualize its explanations according to the audience, ensuring the generated output is tailored to the user’s expertise level. For instance, as noted by [43], persona-based interaction allows the LLM to adjust the technicality of the explanation based on the user’s role, such as a clinician or a system administrator. The prompt informs the LLM in the example that it is “interacting” with a professional clinician, thereby guiding it to generate explanations using clinical terminology and domain-specific knowledge. This capability is critical in clinical settings, where precision and appropriate medical jargon are necessary to convey the physiological implications of the detected anomalies. Moreover, LLM-based explanations can support multi-level decision-making by presenting the expected physiological norms alongside detected anomalies. This dual approach ensures that stakeholders are alerted to deviations and informed of the baseline expectations, facilitating a deeper understanding of abnormal and normal readings. The illustrative example in Figure 2 demonstrates this by asking the LLM to explain why the predicted anomalies occur while requesting the expected physiological behavior when anomaly scores exceed a specified threshold. Here, the system is calibrated to interpret HR and HRV signals, providing reasoning for abnormal readings and expected “normal” physiological patterns using medical language appropriate for the clinical domain.
4 Experiments and Results
@@ 4.1 Experimental Setup
Notice that since the works in emotion monitoring do not model stress detection as an anomaly detection problem (see Table 1), we compare only with the works surveyed in subsection 2.2. We compare supervised UniTS and fine-tuned UniTS with various SoTA models for multivariate time-series anomaly detection. The models compared include TranAD [41], LSTM-NDT [22], CAE-M [46], DAGMM [49], USAD [5], OmniAnomaly [38], MTAD-GAT [47], GDN [12], MAD-GAN [28], MSCRED [44], HypAD [13], and TadGAN [15]. We adopted the hyperparameters presented in their original papers for these models.
Training UniTS
For fine-tuning the UniTS model, we utilized a window size of 5, an embedding dimension of 128, and a learning rate of $5\times 10^{-4}$. The datasets were split into 80% training and 20% testing sets. All models were trained for 20 epochs, and the F1 score was calculated as the average across a 5-fold cross-validation.
Fair comparisons and threshold settings
@@
We systematically preprocessed our datasets to ensure consistency and comparability across different sources. From all datasets, we extracted HR and HRV every 10 seconds using a sliding window approach based on the preceding 60 seconds of signal data from ECG and, when available, Blood Volume Pulse (BVP), which measures the
Prompt:
You are an automated system that interprets changes in HR (Heart Rate) and HRV (Heart Rate Variability) signals when emotional anomalies are detected in a patient’s vital signs with respect to a resting state condition.
The data provided includes:
1. Time: Seconds since the start of the recording 2. HR: Heart Rate in beats per minute 3. HRV: Heart Rate Variability (RMSSD) in milliseconds 4. Prediction: Indicates if an anomaly was detected (1) or not (0) 5. Anomaly Energy: Ratio of the detected anomaly energy to the threshold, indicating the intensity of the anomaly. If energy is $<1$, the sample is normal; otherwise, it indicates an anomaly. The patient is a young adult with no known medical conditions. The data was collected using a wearable ECG sensor.
Patient Data:
Time HR HRV Prediction Anomaly Energy 0 82.39 20.9 0 0.0 80 82.56 10.8 1 1.3
Provide a concise explanation structured in four sections: Analysis, Causes, Criticality, and Scores.
1. Analysis: Highlight any strong, high-intensity anomalies if present, using a dotted list to describe the observed changes in HR and HRV signals. Avoid directly mentioning energy values; instead, describe the anomalies’ clinical relevance and intensity. Describe any trends in HR and HRV over time, noting persistent patterns or shifts in the data that may indicate physiological or emotional responses. 2. Causes: Propose likely causes of these anomalies, considering the patient’s demographic, the characteristics of the detected anomalies (e.g., short, sudden, prolonged), and any environmental or measurement factors that could affect signal accuracy. Distinguish between potential true physiological changes, system errors, or data noise. 3. Criticality: Assess the clinical significance of these anomalies. Consider both the impact of individual anomalies (especially high-energy ones) and the cumulative effect if anomalies are frequent or prolonged. Specify if the anomaly could require immediate follow-up, if it’s non-critical but worth monitoring, or if it seems likely to be insignificant or an error. 4. Scores: (a) Significance Score (0-10): Rate the significance of the anomaly based on how clinically relevant it is, with 0 being not relevant (e.g., noise) and 10 being highly relevant (e.g., real physiological change). (b) Criticality Score (0-10): Rate the potential clinical impact or urgency, with 0 being minimal concern and 10 being critical. This should consider both the intensity of the anomaly and the risk it may pose to the patient’s well-being. Ensure the explanation is brief, clinically relevant, and suitable for interpretation by a professional clinician.
Figure 2: Prompt used to interpret HR and HRV signals, focusing on emotional anomalies, and generate a structured clinical analysis for patient data interpretation.
volumetric change in blood circulation through the microvascular bed of tissue. BVP signals can be extracted non-invasively using photoplethysmography (PPG), a technique commonly integrated into smartwatches and other wearable devices. This capability is instrumental in stress detection and management, making wearable devices valuable in clinical and everyday settings. Lastly, we use a threshold of the 3rd and 97th percentile to indicate anomalous data points: i.e., everything below the 3rd or above the 97th percentile is labeled as 1; otherwise, it is 0. Percentiles are calculated on the reconstruction errors.
@@ 4.2 Datasets
We use the following three widely adopted datasets for emotion detection:
  • 1.
    DREAMER [23] is designed for emotion recognition via low-cost devices. It consists of EEG and ECG readings recorded from 23 participants. ECG signals were captured at a sampling rate of 256 Hz using a SHIMMER wireless sensor. Movie clips were used to stimulate different emotions in participants, and at the end of each clip, participants scored their self-assessment on a 5-point arousal, valence, and dominance scale. Here, we treat states with a valence of 1 as abnormal and the rest as normal.
  • 2.
    MAHNOB-HCI [37] was collected using the Biosemi Active II system with active electrodes to capture physiological signals, including ECG and EEG, from 27 participants at a sampling rate of 256 Hz. Participants were asked to record their emotional status after each trial, consisting of multimedia content stimulation. Here, we treat the states with a valence of 1 as abnormal and the rest of the emotional states as normal. For our experiments, we selected subsets of the dataset with recordings of at least 120 seconds and excluded those with incomplete tracks due to recording errors.
  • 3.
    WESAD [36] has ECG and BVP signals from 15 participants at a rate of 700Hz extracted via a RespiBAN Professional sensor worn on the chest and the Empatica E4 wristband. It contains four different emotional states. Here, we treat the stress state as abnormal and the rest as normal. The stress condition is obtained by exposing the subjects to the Tier Social Stress Test, which consists of public speaking and a mental arithmetic task.
Note that all datasets induce emotions in a controlled setting, a limitation that we discuss in subsection 4.4.
Table 3: Comparison of different HR and HRV methods extracted from ECG and BVP. Bold-faced values illustrate the best; underlined ones are the second best.
@@
LSTM-NDT [22] 0.235 0.373 0.694 0.760 0.515 ± 0.218
MSCRED [44] 0.434 0.618 0.594 0.536 0.545 ± 0.071
HypAD [13] 0.513 0.547 0.669 0.479 0.552 ± 0.072
TranAD [41] 0.515 0.622 0.474 0.717 0.582 ± 0.095
OmniAnomaly [38] 0.543 0.513 0.716 0.583 0.589 ± 0.078
MTAD-GAT [47] 0.369 0.677 0.757 0.676 0.620 ± 0.149
CAE-M [46] 0.634 0.513 0.762 0.590 0.625 ± 0.090
MAD-GAN [28] 0.390 0.678 0.735 0.750 0.638 ± 0.146
TadGAN [15] 0.528 0.657 0.705 0.680 0.643 ± 0.068
USAD [5] 0.590 0.641 0.686 0.759 0.669 ± 0.062
GDN [12] 0.712 0.501 0.697 0.778 0.672 ± 0.103
DAGMM [49] 0.619 0.596 0.744 0.733 0.673 ± 0.066
UniTS [14] 0.828 0.864 0.793 0.800 0.821 ± 0.049
Figure 3: Confidence intervals for the performances of UniTS and the best SoTA methods as per the Dunn post-hoc test. We highlight the method that performs best on average according to the values reported in Table 3.
@@ 4.3 Quantitative Results
UniTS surpasses SoTA anomaly detection systems and shows remarkable stability and robustness across datasets
@@
For completeness purposes, in Figure 3, we provide the reader with confidence intervals on the performances of UniTS and five top-performing SoTA methods according to the Dunn post-hoc test analysis – see Table LABEL:tab:dunn_p_values.
UniTS achieves performances from lightweight devices comparable to SoTA systems using more invasive and potentially accurate sensors
The only dataset collecting signals from ECG devices and wristbands (BVP) is WESAD. As shown in Table 3, UniTS delivers consistently high F1 scores with both ECG and BVP and maintains similar results across both types, while some other systems experience varied performance with BVP signals. This is a valuable result since, as discussed in section 1, lightweight devices, such as wristbands and smartwatches, are less expensive and invasive, enabling seamless monitoring. Using non-invasive, wearable technology without sacrificing performance can expand stress-related research and applications. Future studies can integrate these devices in various settings to gather continuous stress data, leading to personalized stress management programs, early intervention systems for mental health, and better well-being monitoring.
[…]
Figure 5 shows the importance of the sampling rate on all datasets on ECG and BVP versions. Here, we illustrate the trend of average F1 scores of UniTS when the sampling interval increases from 10 to 60 seconds. The vital signals (i.e., HR and HRV) are sampled every $k$ seconds to build the multivariate time series taken in input, with varying $k$. As expected, the longer the sampling interval, the less responsive the model becomes, resulting in a degradation of results. This phenomenon is even more visible when we extract the HR and HRV from BVP, having a drop of $-53.89\%$ passing from a 10 to a 60-second sampling rate. This discrepancy highlights a critical consideration in the evolving landscape of wearable technology and continuous monitoring systems. Therefore, we suggest re-evaluating the balance between device convenience and performance. For applications requiring high precision, such as medical diagnostics or performance monitoring in athletes, adhering to more frequent sampling intervals appears imperative. Therefore, our study advocates for a nuanced approach, where the selection of sampling intervals is tailored to the specific requirements of the use case, ensuring that the constraints of the monitoring device do not compromise the integrity of physiological data.
Figure 5: Performance degradation of F1 score across different sampling intervals using the UniTS on all the datasets. The zoomed-in area illustrates how the overall decreasing performance trend starts to have diminishing effects, especially in WESAD (BVP).
4.4 Qualitative Results
@@
Casting the problem as one of anomaly detection enhances explainability
Anomaly detection begins by establishing a model of “normal” behavior based on the monitored signals. Next, it detects deviations from this normality. This helps clinicians understand the baseline deviations identified and align better with medical professionals’ needs and practical constraints. First, clinicians are usually more interested in identifying and investigating abnormalities than confirming the norm. Anomaly detection systems align with doctors’ diagnostic workflow by focusing on detecting deviations from the norm. Secondly, these models provide significant benefits for precision medicine and personalization due to their ability to identify unique patient profiles and adapt to new data continuously [33]. Furthermore, in the case of multiple monitored signals, the model should highlight which ones are more indicative of stress [4]. In detail, although anomaly detectors are not interpretable per se, [^10] Without loss of generality, anomaly detectors produce reconstruction errors for each data point, combining all feature dimensions. These errors do not necessarily indicate which features/signals contribute to the anomalousness of the data point itself. UniTS reconstructs the multivariate time series to provide reconstruction errors for each series’ dimension. This inherent characteristic allows us to interpret whether anomalies occurred, which is paramount, and which features contributed to their detection.
@@
For example, Figure 6 shows how to support doctors in identifying which signals contributed the most to detecting anomalies. Here, we depict the contribution probability based on the reconstruction error of each signal – i.e., HR (up) and HRV (middle) – in determining anomalies. The more opaque the shade of each bar in the plot, the more likely an anomaly will occur. Notice how the opacity of the probabilities increases when there is an abrupt trend change in one (or both) of the signals, which aligns with the red areas depicting the detection of an anomaly. One can think of the transparency of each bar as a visual uncertainty of the model. Therefore, the more opaque the bars are, the more certain the model is that there is an anomaly in that particular time frame. Meanwhile, when the bars are transparent, they may indicate two things: either the model is uncertain, or the data points therein are normal. To this end, we integrate the contribution probability with detecting anomalies in the raw signals. Additionally, we invite the reader to notice the different transparency levels within the time frame where an anomaly is detected. This phenomenon is expected as not all anomalies (gradual vs. abrupt) are detected with the same confidence. Interestingly, badly reconstructing HR in this scenario pushes the model to be more confident about an anomaly occurring – see the middle portion of the third row in Figure 6. Contrarily, even if UniTS cannot effectively reconstruct HRV, it does not lead to detecting anomalies – e.g., notice the more transparent tail-ends of the probabilities plot. By exploiting these types of explicit and inherently interpretable visualizations – i.e., the contribution probability is estimated via the reconstruction error – doctors can interpret which signal guides the prediction of an anomaly during inference and intervene accordingly. Finally, this type of analysis can be used to debug if the monitoring system functions correctly. For instance, if the contribution probability bars are opaque, yet the system does not report an anomaly, there might be something wrong with the current monitoring setup. In addition to a graphical representation of the anomaly (showing the dimensions that contribute most to the detected anomaly), we provide clinicians with a rich textual explanation, as detailed below.
Zero-shot integration with LLMs further enhances explainability
In real-world scenarios, anomalies detected by the model require further investigation by a medical professional to ensure their correctness and assess whether they could result in any significant health risks for the patient. The anomaly detection model offers useful insights, but these findings require more context and interpretation. Using an LLM, we can provide detailed explanations that help clarify the anomaly and its potential impact, as illustrated in Figure 9 and Figure 10 of C. Here, we evaluated the utility and accuracy of explanations generated by GPT-4o using qualitative assessments from a medical professional. The explanations, derived from the outputs of the anomaly detection model, were structured into four sections: Analysis, Causes, Criticality, and Scores, according to the scheme illustrated in Figure 2. We generated an evaluation set of 16 explanations, which were reviewed by a highly experienced doctor, [^11] With more than 40 years of experience. who thoroughly examined the clinical data and images before considering the textual explanations provided by the system. Based on her evaluation, we assessed the explanations’ relevance, coherence, and specificity, with the results summarized in Table 5.
Table 5: Qualitative evaluation by a medical professional of GPT-4o’s explanations, assessing the usefulness of each analysis, causes, criticality, and the coherence and specificity of the overall explanation. Ratings are on a 0-5 scale. Each dataset contains 4 different explanations. Here, we show averages for each dataset and the cumulative average ($\mu\pm\sigma$) over all explanation evaluations.
Utility of
Analysis
Utility of
Causes
Utility of
Criticality
Explanation
Coherence
Explanation
Specificity
DREAMER 5.00 4.75 4.50 4.25 1.25
HCI 5.00 4.75 4.50 4.50 0.50
WESAD (ECG) 4.75 4.25 4.50 4.50 0.25
WESAD (BVP) 4.75 4.75 5.00 4.75 0.75
$\mu\pm\sigma$ $4.875\pm 0.331$ $4.625\pm 0.484$ $4.625\pm 0.599$ $4.500\pm 0.500$ $0.688\pm 1.261$
The results indicate that GPT-4o provided highly useful explanations, with average utility scores of 4.875 for analysis, 4.625 for causes, and 4.625 for criticality. While the overall coherence of the explanations was strong, with an average score of 4.5, the low specificity score of 0.688 can be attributed to two main factors. First, as we already remarked, the datasets were collected in a controlled setting, which provided limited contextual information about the patient’s activities or other physiological signals. This restricted the ability to contextualize the anomalies with a broader understanding of the patient’s behavior or lifestyle. Second, the lack of detailed clinical history for the patient further limited the depth of the explanations. Without access to relevant background information, the model could only offer general insights based on the detected anomalies rather than providing a more personalized interpretation tied to the patient’s specific condition or medical history. The medical professional was also asked to rate the significance of the anomaly to assess its relevance, as well as its criticality to evaluate the potential risk it poses to the patient, taking into account factors such as age and overall health condition. These ratings were then compared with those provided by GPT-4o to determine whether numerical metrics could be generated to offer an immediately comprehensible score. Such a score would help clinicians rapidly identify anomalies that are of particular concern and may require urgent attention or further investigation. It should be noted that while GPT-4o generated ratings on a 0-10 scale (see Figure 2), the medical professional used a more compact 0-5 scale. To use a finer-grained scale to reduce the effect of small rating errors, we requested GPT-4o to produce results on a 0-10 scale. In this way, we obtain step errors leading to a smaller fractional error when mapped to the 0-5 scale originally used by the doctor. Afterward, the GPT-4o ratings were rescaled by dividing by 2, aligning them to the 0-5 range for direct comparison with the doctor’s ratings.
The results are presented in Table 6 – for error distributions, refer to Figure 7 – which shows that GPT-4o generally assigned higher ratings across all metrics and datasets. However, the average error rates were 0.938 for significance and 0.688 for criticality, indicating that the model’s assessments were comparable to the expert’s. These findings suggest that by using an LLM, we can generate useful explanations that support and accelerate the review process of anomalies, providing valuable assistance to medical professionals in assessing and prioritizing potential issues.
Table 6: Comparison of anomaly significance and criticality ratings between a medical professional and GPT-4o, showing the errors in GPT estimations for each dataset. The ratings are on a 0-5 scale. Each dataset contains 4 explanations. Here, we show averages for each dataset and the cumulative average ($\mu\pm\sigma$) over all explanation evaluations.
Significance Criticality
Human
Evaluation
GPT
Evaluation
Error
Human
Evaluation
GPT
Evaluation
Error
DREAMER 1.00 2.25 1.25 0.25 1.25 1.00
HCI 1.00 2.13 1.13 0.50 1.00 0.50
WESAD (ECG) 1.50 2.25 0.75 0.50 1.50 1.00
WESAD (BVP) 1.50 2.13 0.63 0.75 1.00 0.25
$\mu\pm\sigma$ $0.938\pm 0.496$ $0.688\pm 0.496$
Figure 7: Distribution of errors in GPT-4o’s anomaly significance (left) and criticality ratings (right) compared to a medical professional’s assessments. Ratings are on a scale from 0 to 5.
5 Discussion and Conclusion
Summary of results
This is the first study to employ a universal time series model, UniTS, to support the seamless monitoring of physiological signals. We conducted an extensive evaluation by benchmarking UniTS against 12 leading methods across three datasets for detecting stress and abnormal emotions. Our results show that UniTS consistently outperforms state-of-the-art models (see subsection 4.3). Notably, UniTS performs equally well when using both invasive monitoring devices, such as ECG, and more lightweight alternatives like wristbands, demonstrating strong resilience to noisy data. This study also examines how the sampling interval influences the model’s performance, demonstrating that robust results can be achieved even with lower-cost monitoring equipment as long as the sampling interval stays within acceptable thresholds.
Benefits in clinical applications
In the literature, a combination of physiological, behavioral, and psychological monitoring methods have been proposed to monitor the stress level in patients with neurodegenerative diseases, facilitated by advanced wearable devices and digital technologies  [39]. Speech and video analysis, ECG, EEG, and implantable devices have shown significant promise in stress monitoring for patients with neurodegenerative diseases. These methods have been shown to provide more accurate, continuous, and detailed insights into stress levels than less invasive approaches. However, they present challenges such as higher costs, potential discomfort, privacy concerns, and the need for manual procedures. In this paper, we demonstrated that using last-generation foundational AI models can boost the performance of less invasive devices, making them comparable to those of devices with a greater impact on the daily lives of monitored patients. Monitoring the progression of patients’ disease using lightweight devices has numerous advantages in clinical practice, including increased patient comfort and compliance, continuous and real-time monitoring, reduced healthcare costs, increased patient independence, personalized care, improved remote patient management, and improved health outcomes through early intervention.
Furthermore, we have shown how casting the stress detection problem as one of anomaly detection, rather than classification, improves the explainability of the model’s outcomes (subsection 4.4). Explainability is vital in medical applications of AI because it ensures trust and accountability, supports clinical validation, and facilitates integration into clinical workflows. In addition, anomaly detection better aligns with the clinical practice since each patient has unique physiological and behavioral baselines. Anomaly detection allows for creating personalized baselines, making identifying deviations that indicate stress easier.
Limitations and future work